A company sells equipment for $6,000. The original cost was $50,000. The accumulated depreciation is $45,000. The sale results in


A. a loss of $1,000


B. a gain of $1,000


C. a gain of $5,000


D. neither a gain nor a loss
Answer #1

Answer: 


Option B "a gain of $1,000" 


Explanation 


Book value = 50000 - 45000 = 5000 


Gain on sale = 6000 - 5000 = 1000

answered by: Ashik Padanilam